# GravSphere2: A higher-order Jeans method for mass-modeling spherical stellar systems

### v1

@@ Appendix A Velocity distribution function
Given the methodology discussed in Sec. 2 to derive velocity moments from a given mass model, we now need a prescription to constrain these moments from observable velocities. A common approach is to bin the data into discrete portions of a certain spatial width. While this has been shown to be sufficient to well-recover mass models from dwarf galaxy mock data (Read & Steger 2017; Read et al. 2018; Gregory et al. 2019; Genina et al. 2020; Read et al. 2021; Collins et al. 2021; De Leo et al. 2024; Nguyen et al. 2025; Tchiorniy & Genina 2025), it can lead to a loss of information and potential biases. It introduces an unavoidable trade-off between spatial and velocity resolution, which may be particularly acute when aiming to extract information of higher-order moments, and it makes it challenging to properly marginalize over uncertain membership for the tracers, survey selection effects, binary stars, etc.
@@
To solve all of the above problems, in GravSphere2 we model each stellar velocity individually, without binning. In the context of Jeans analyses, comparable approaches have been introduced in the past, such as DiscreteJAM (Cappellari 2008; Watkins et al. 2013; Cappellari 2015), MAMPOSSt (Mamon et al. 2013), and more recently in Wardana et al. (2025) [^24] The approach from Wardana et al. (2025), which was recently and independently developed, is the one most similar to the one we use in GravSphere2. Some of the main differences between these include the improvements which we introduce, such as a fully general treatment of anisotropies (we marginalize $\beta$ and $\beta^{\prime}$ as two independent and radially dependent, i.e. non-constant, quantities) and the use of PMs to constrain both the second and fourth-order Jeans equations. We also implement uncertainties differently in our PDF modeling, which may explain why these authors find greater susceptibility to error-driven biases in their results. . In our case, rather than assuming Gaussianity for the velocity PDF (projected or otherwise), we use the PDF models introduced in Sanders & Evans (2020). These are based on analytical convolutions of a Gaussian with a uniform distribution kernel for models of negative excess kurtosis and with a Laplace kernel for positive excess kurtosis models. Together, these allow meaningful deviations from Gaussianity with kurtoses $\kappa\equiv\braket{v^{4}}/\sigma^{4}$ in the range $1.8<\kappa<6$, with $\kappa=3$ corresponding to a Gaussian (which defines excess kurtosis). For a symmetric distribution of mean zero for the scaled velocity variable $w\equiv v/\sigma$ and measurement error $s\equiv\delta v/\sigma$ (assumed to be symmetric and Gaussian), this yields the error-convolved PDF of form:
@@
and the parameters $a$ and $b$ are fully specified by the variance and kurtosis (respectively) as: [^25] Note that this is the intrinsic variance and kurtosis of the velocity distribution before the error convolution.
@@
This allows for a fully self-consistent treatment between the assumed PDF and the stellar kinematic model. [^26] Our numerical implementation of these functions is based on the code introduced in Sanders & Evans (2020), which is publicly available at: https://github.com/jls713/gh_alternative.
Appendix B Mass-anisotropy modeling
As in Bañares-Hernández et al. (2025), for the photometric tracer density profile, we use the generalized $\alpha\beta\gamma$ profile (e.g. Hernquist 1990; Zhao 1996). This is given by a double power-law model:
$\nu_{\star}(r)=\frac{\rho_{c}}{(r/r_{c})^{\gamma}(1+(r/r_{c})^{\alpha})^{(\beta-\gamma)/\alpha}},$ (33)
where we have introduced the three exponent variables $\alpha,\beta,\gamma,$ and the scale radius and density $r_{c},\rho_{c},$ respectively.
Following recent versions of GravSphere (e.g. Collins et al. 2021), we use the coreNFWtides profile (Read et al. 2018) to model dark matter profiles of tidally stripped galactic halos expected from hydrodynamical simulations simulating Local Group environments. [^27] Previous versions of GravSphere  have used non-parametric modeling with sequential power laws Read & Steger (2017); Genina et al. (2020). As compared to these earlier choices, the coreNFWtides profile has number of advantages: it does not suffer from biases due to monotonicity priors inherent to the previous method (Tchiorniy & Genina 2025), it is designed to model mass profiles realistic galactic systems, and it makes recovering cosmologically valuable parameters such as the halo mass and concentration straightforward. This has an enclosed mass profile given by:
@@ Appendix B Mass-anisotropy modeling
is the coreNFW profile (Read et al. 2016) and we have the original Navarro-Frenk-White (NFW) profile (Navarro et al. 1996):
@@
with $c_{200}$ being the dimensionless concentration parameter, and $r_{200},\>M_{200}$ being the virial radius and mass (respectively), defined at a region where the average density exceeds the present-day cosmological critical density $\rho_{\rm crit}=135.05\>{\rm M}_{\odot}\>{\rm kpc}^{-3}$ by the critical overdensity factor $\Delta=200$. [^28] The precise value of the adopted cosmological critical density will vary depending on the adopted cosmology. Our value corresponds to the ‘canonical value’ with Hubble parameter $H_{0}\sim 70$ km s-1/Mpc, so that $\rho_{c}=3H_{0}^{2}/8\pi G$. It should also be noted that different authors use different values of $\Delta$ to define the virial radius. The core formation parameter $n$ yields full core formation when $n=1,$ retrieves an NFW cusp when $n=0,$ and can model potentially cuspier distributions for negative values. The core radius $r_{c}$ determines the extent of the inner core region. The tidal radius $r_{t}$ and slope exponent $\delta$ regulate the amount of stripping in the outer region.
Following previous versions of GravSphere, we use the anisotropy profile assuming the following the generalized form from Baes & van Hese (2007):
@@
where $\beta_{0}$ is the anisotropy at the center, $\beta_{\infty}$ its limit as it asymptotically approaches infinity, $r_{\beta}$ is the transition scale between these two regimes, and $\eta$ is the exponent which modulates the steepness of the transition. We also assume the same functional form for the fourth-order anisotropy $\beta^{\prime},$ allowing for a fully general treatment of this quantity (unlike previous works that have made more restrictive assumptions about its form).
Appendix C Likelihood and priors
@@
where $\boldsymbol{\theta}$ includes all the model parameters, which are simultaneously fitted. This includes the anisotropy parameters (that are distinct for the second and fourth order anisotropy profiles), the mass model parameters, and the photometric tracer profile parameters. The first three terms correspond to the likelihood due to individual velocity PDFs (Eqs. (25) and (A)) for the LOS and tangential and radial PM components for the data we consider, whilst the last term corresponds to the term from the photometric tracer profile. GravSphere2 incorporates two options in this respect: one for a binned photometric light profile with Gaussian errors, and a completely ‘bin-free’ option where individual stellar tracer positions are used instead. [^29] Note that GravSphere2 does not bin any data. For the case of the binned profile it is assumed that this is provided as a data constraint from observations. Note that it would not be necessary to derive a binned profile from individual stars (unlike previous versions of GravSphere) as one could then simply use the ‘bin-free’ approach instead. In the latter case, much like with the kinematic tracers, one can write the total likelihood as the product of those of individual photometric tracers, whose (positional) PDFs take the form:
@@
In all cases, we verified that our fits were able to reproduce the photometric tracer profiles well and recover the true solution, which is easily achievable with the general $\alpha\beta\gamma$ functional form (Eq. 33). [^30] We have omitted showing these results since Plummer-like distributions such as the ones from our mock sample are quite trivial to reproduce, and also due to the fact that we did not always use binned data. The advantages of the $\alpha\beta\gamma$ functional form will be more evident when reproducing more complex distributions which cannot be reduced to Plummer profiles (or a summation of these), this includes cuspier distributions or those with shallow slopes at higher radii. One example of this is the GC Omega Centauri, where this functional form was introduced for GravSphere(Bañares-Hernández et al. 2025).
As in Bañares-Hernández et al. (2025), we work with the symmetrized anisotropy (for both the second and fourth order anisotropy profiles):
@@
For the Gaia Challenge data (Sec. 3.1), the priors we adopt for the mass and anisotropy profile are the same as those in Section 4.1.2 of Collins et al. (2021), whilst for the photometric component we allow a flat prior with limits set with 50% scatter around the true solution, which we fit as a binned surface density profile using 10,000 tracers.
@@
For the Fornax-like simulated dwarf galaxy (Sec. 3.2), our mass and anisotropy priors are also the same as those from Collins et al. (2021) except for the fact the we use unbounded values for the limiting anisotropies ($1<\widetilde{\beta}_{0,\infty}<1$) and broader priors for the concentration parameter ($1<c_{200}<50$). For the stellar tracer model, we do not use binned photometry and instead use the full ‘bin-free’ method, using Eq. (45). We allow a scatter of 1 dex from the 2D half-light radius ($R_{1/2}$) of $\sim 700$ pc for the length scale in the photometric light profile (which we sample logarithmically as a flat prior) [^31] This can be obtained directly from the photometric tracers (either from the profile or their median position), but is only meant to set a relevant scale and its precise value is not needed as the results are insensitive to it. and $0.1<\alpha<4,\>3.1<\beta<7,\>0<\gamma<2.9$. [^32] These are chosen to allow for convergence (needed to set a proper normalization) while allowing a wide range of profiles of varying steepness. (Recall that $\gamma$ and $\beta$ are the negative inner and outer logarithmic slopes of the light profile (respectively), where a Plummer profile corresponds to $\alpha,\>\beta,\>\gamma=2,5,0).$ Lastly, we sample the mass of the stellar component with a flat prior with 25% scatter around $4.3\times 10^{7}\>\text{M}_{\odot},$ as in Tchiorniy & Genina (2025), assuming a constant mass-to-light ratio and that the stellar mass distribution has the same radial density profile as the tracer stars. [^33] This is usually a good approximation provided that the tracer profile is complete and that there is no meaningful mass segregation or observational bias in the photometric selection. This is typically a good approximation for nearby dwarf galaxies, which have relaxation times longer than a Hubble time due to the presence of dark matter (e.g. Baumgardt et al. 2022). If this is not the case, as in star clusters, one needs to include the presence of additional components of varying segregation (e.g. see Dickson et al. 2023; Bañares-Hernández et al. 2025) in the mass model.
@@
The full implementation of our likelihood modeling with the data we use, and the GravSphere2 code itself, is made publicly available in the Introduction (Sec. 1).
@@
All fits were performed using the nested sampling package dynesty (Speagle 2020; Koposov et al. 2022). This is based on nested sampling techniques (Skilling 2004, 2006) and the ‘dynamic nested sampling technique’ (Higson et al. 2019) (which we use here). dynesty has been optimized for the inference of posterior distributions (as opposed to evidence estimation), using the bounding method from Feroz et al. (2009). The sampling method used is based on Neal (2003); Handley et al. (2015a, b). [^34] For further information, see: https://dynesty.readthedocs.io
/en/latest/index.html
. An example with the specific implementation used for GravSphere2 can be found in the publicly available code release.
@@
This sampling method has been shown to outperform Markov-chain Monte Carlo Methods (MCMC), particularly when dealing with complex, high dimensional, and multi-modal posterior spaces (e.g. Speagle 2020). This is consistent with our own experiments comparing with previous versions of GravSphere that used emcee instead (Foreman-Mackey et al. 2013). Following Bañares-Hernández et al. (2025), we have thus opted to use this improved sampling technique for GravSphere2.

### v2

@@ 2.2 Velocity distribution function
Given the methodology discussed in Sec. 2.1 to derive velocity moments from a given mass model, we now need a prescription to constrain these moments from observable velocities. A common approach is to bin the data into discrete portions of a certain spatial width. While this has been shown to be sufficient to well-recover mass models from dwarf galaxy mock data (2017MNRAS.471.4541R; 2018MNRAS.481..860R; 2019MNRAS.485.2010G; 2020MNRAS.498..144G; 2021MNRAS.501..978R; 2021MNRAS.505.5686C; 2024MNRAS.535.1015D; 2025arXiv250303812N; 2025arXiv250418617T), it can lead to a loss of information and potential biases. It introduces an unavoidable trade-off between spatial and velocity resolution, which may be particularly acute when aiming to extract information of higher-order moments, and it makes it challenging to properly marginalize over uncertain membership for the tracers, survey selection effects, binary stars, etc.
@@
To solve all of the above problems, in GravSphere2 we model each stellar velocity individually, without binning. In the context of Jeans analyses, comparable approaches have been introduced in the past, such as DiscreteJAM (2008MNRAS.390...71C; 2013MNRAS.436.2598W; 2015arXiv150405533C), MAMPOSSt (2013MNRAS.429.3079M), and more recently in 2024arXiv240412671W [^9] The approach from 2024arXiv240412671W, which was recently and independently developed, is the one most similar to the one we use in GravSphere2. Some of the main differences between these include the improvements which we introduce, such as a fully general treatment of anisotropies (we marginalize $\beta$ and $\beta^{\prime}$ as two independent and radially dependent, i.e. non-constant, quantities) and the use of PMs to constrain both the second and fourth-order Jeans equations. We also implement uncertainties differently in our PDF modeling, which may explain why these authors find greater susceptibility to error-driven biases in their results. . In our case, rather than assuming Gaussianity for the velocity PDF (projected or otherwise), we use the PDF models introduced in 2020MNRAS.499.5806S. These are based on analytical convolutions of a Gaussian with a uniform distribution kernel for models of negative excess kurtosis and with a Laplace kernel for positive excess kurtosis models. Together, these allow meaningful deviations from Gaussianity with kurtoses $\kappa\equiv\braket{v^{4}}/\sigma^{4}$ in the range $1.8<\kappa<6$, with $\kappa=3$ corresponding to a Gaussian (which defines excess kurtosis). For a symmetric distribution of mean zero for the scaled velocity variable $w\equiv v/\sigma$ and measurement error $s\equiv\delta v/\sigma$ (assumed to be symmetric and Gaussian), this yields the error-convolved PDF of form:
@@
and the parameters $a$ and $b$ are fully specified by the variance and kurtosis (before convolving with errors), respectively, as:
@@
This allows for a fully self-consistent treatment between the assumed PDF and the stellar kinematic model. [^10] Our numerical implementation of these functions is based on the code introduced in 2020MNRAS.499.5806S, which is publicly available at: https://github.com/jls713/gh_alternative.
2.3 Mass-anisotropy modeling
As in 2025A&A...693A.104B, for the photometric tracer density profile, we use the generalized $\alpha\beta\gamma$ profile (e.g. 1990ApJ...356..359H; 1996MNRAS.278..488Z). This is given by a double power-law model:
$\nu_{\star}(r)=\frac{\rho_{p}}{(r/r_{p})^{\gamma}(1+(r/r_{p})^{\alpha})^{(\beta-\gamma)/\alpha}},$ (33)
where we have introduced the three exponent variables $\alpha,\beta,\gamma,$ and the scale radius and density $r_{p},\rho_{p},$ respectively.
Following recent versions of GravSphere (e.g. 2021MNRAS.505.5686C), we use the coreNFWtides profile (2018MNRAS.481..860R) to model dark matter profiles of tidally stripped galactic halos expected from hydrodynamical simulations simulating Local Group environments. Previous versions of GravSphere  have used non-parametric modeling with sequential power laws 2017MNRAS.471.4541R; 2020MNRAS.498..144G. As compared to these earlier choices, the coreNFWtides profile has number of advantages: it does not suffer from biases due to monotonicity priors inherent to the previous method (2025arXiv250418617T), it is designed to model mass profiles realistic galactic systems, and it makes recovering cosmologically valuable parameters such as the halo mass and concentration straightforward. The model has an enclosed mass profile given by:
@@ 2.3 Mass-anisotropy modeling
is the coreNFW profile (Read:2015sta) and we have the original Navarro-Frenk-White (NFW) profile (1996ApJ...462..563N):
@@
with $c_{200}$ being the dimensionless concentration parameter, and $r_{200},\>M_{200}$ being the virial radius and mass (respectively), defined at a region where the average density exceeds the present-day cosmological critical density $\rho_{\rm crit}=135.05\>{\rm M}_{\odot}\>{\rm kpc}^{-3}$ by the critical overdensity factor $\Delta=200$. [^11] The precise value of the adopted cosmological critical density will vary depending on the adopted cosmology. Our value corresponds to the ‘canonical value’ with Hubble parameter $H_{0}\sim 70$ km s-1/Mpc, so that $\rho_{\rm crit}=3H_{0}^{2}/8\pi G$. It should also be noted that different authors use different values of $\Delta$ to define the virial radius. The core formation parameter $n$ yields full core formation when $n=1,$ retrieves an NFW cusp when $n=0,$ and can model potentially cuspier distributions for negative values. The core radius $r_{c}$ determines the extent of the inner core region. The tidal radius $r_{t}$ and slope exponent $\delta$ regulate the amount of stripping in the outer region.
Following previous versions of GravSphere, we use the anisotropy profile assuming the following the generalized form from 2007A&A...471..419B:
@@
where $\beta_{0}$ is the anisotropy at the center, $\beta_{\infty}$ its limit as it asymptotically approaches infinity, $r_{\beta}$ is the transition scale between these two regimes, and $\eta$ is the exponent which modulates the steepness of the transition. We also assume the same functional form for the fourth-order anisotropy $\beta^{\prime},$ allowing for a fully general treatment of this quantity, unlike previous works that have made more restrictive assumptions about its form (e.g. 1990AJ.....99.1548M; 2003MNRAS.343..401L; 2005MNRAS.363..918L; 2013MNRAS.432.3361R; 2024arXiv240412671W).
3 Data, likelihood, and priors
3.1 Gaia Challenge mock data
To test the overall performance of GravSphere2 we first use a subset of the Gaia Challenge catalog of mock galaxies [^12] See the Gaia Challenge website for full access to the data and further information: https://astrowiki.surrey.ac.uk/., constructed from spherical dynamical equilibrium models. The results are shown in Sec. 4.1. These have the advantage of having been extensively used for prior testing in the literature, allowing for a straightforward comparison with previous studies, including previous versions of GravSphere (e.g. 2019MNRAS.485.2010G; 2021MNRAS.501..978R; 2021MNRAS.505.5686C). As in 2021MNRAS.505.5686C, we focus on a subset of two mock galaxies originally from 2011ApJ...742...20W which have been the ones for which several previous methods (including GravSphere) have been affected by significant bias (e.g. see 2021MNRAS.501..978R). These correspond to a dark matter profile with the $\alpha\beta\gamma$ functional form (same as Eq. 33) embedded in a Plummer-like [^13] In practice, these profiles closely resemble a Plummer sphere but show minor deviations in the form of an $\alpha\beta\gamma$ profile (Eq. 33), which ensures positivity of the phase-space distribution function. stellar tracer distribution, with a varying Osipkov-Merritt anisotropy profile (1979PAZh....5...77O; 1985MNRAS.214P..25M). These two mock galaxies consist of a cored dark matter model (PlumCoreOM) and a cuspy one (PlumCuspOM). We consider cases with 100, 1,000, and 10,000 tracers, with and without including PMs. In all cases we simultaneously fit the binned surface density profile using 10,000 tracers, assuming also velocities with 2 km/s errors.
3.2 Simulated dwarf galaxy data
The second case we consider is the recent suite of simulated dwarf galaxies from 2025arXiv250418617T. These are initialized with cuspy NFW profiles with massive Plummer-like stellar components and are evolved in a Milky Way-like potential to simulate the properties of nearby present-day dSphs. This allows us to probe more realistic galactic environments and aspects such as departures of equilibrium and departures from spherical symmetry due to tidal effects, or induced rotation, which may be present in simulations.
[…]
3.3 Likelihood and priors
@@
where $\boldsymbol{\theta}$ includes all the model parameters, which are simultaneously fitted. This includes the anisotropy parameters (that are distinct for the second and fourth order anisotropy profiles), the mass model parameters, and the photometric tracer profile parameters. The first three terms correspond to the likelihood due to individual velocity PDFs (Eqs. (25) and (2.2)) for the LOS and tangential and radial PM components for the data we consider, whilst the last term corresponds to the term from the photometric tracer profile. GravSphere2 incorporates two options in this respect: one for a binned photometric light profile with Gaussian errors, and a completely ‘bin-free’ option where individual stellar tracer positions are used instead. Note that GravSphere2 does not bin any data. For the case of the binned profile it is assumed that this is provided as a data constraint from observations. Unlike previous versions of GravSphere, it is not necessary to derive a binned profile from individual stars as one can simply use the ‘bin-free’ approach instead. In the latter case, much like with the kinematic tracers, one can write the total likelihood as the product of those of individual photometric tracers, whose (positional) PDFs take the form:
@@
In all cases, we verified that our fits were able to reproduce the photometric tracer profiles well and recover the true solution. We show a representative example fit in Appendix 9 (Fig. 10), showing excellent recovery which is far in excess of the range allowed by our priors. The advantages of the $\alpha\beta\gamma$ functional form will be more evident when reproducing more complex distributions which cannot be reduced to Plummer profiles (or a summation of these), this includes cuspier distributions or those with shallow slopes at higher radii. One example of this is the GC Omega Centauri, where this functional form was introduced for GravSphere(2025A&A...693A.104B).
As in 2025A&A...693A.104B, we work with the symmetrized anisotropy (for both the second and fourth order anisotropy profiles):
@@
For the Gaia Challenge data (Sec. 3.1), the priors we adopt for the mass and anisotropy profile are similar (albeit broader for some quantities) to those Section 4.1.2 of 2021MNRAS.505.5686C, whilst for the photometric component we allow a flat prior with limits set with 50% scatter around the true solution, which we fit as a binned surface density profile using 10,000 tracers. A summary of all priors adopted is shown in Table 17.
@@
For the Fornax-like simulated dwarf galaxy (Sec. 3.2), our mass and anisotropy priors are also the same as for the Gaia Challenge mocks except for the fact the we use unbounded values for the limiting anisotropies ($1<\widetilde{\beta}_{0,\infty}<1$) and somewhat less broad priors for the concentration parameter ($1<c_{200}<50$). For the stellar tracer model, we do not use binned photometry and instead use the full ‘bin-free’ method, using Eq. (45). We allow a scatter of 1 dex from the 2D half-light radius ($R_{1/2}$) of $\sim 770$ pc for the length scale in the photometric light profile (which we sample logarithmically as a flat prior) [^16] This can be obtained directly from the photometric tracers (either from the profile or their median position), but is only meant to set a relevant scale and its precise value is not needed as the results are insensitive to it. and $0.1<\alpha<4,\>3.1<\beta<7,\>0<\gamma<2.9$. These are chosen to allow for convergence (needed to set a proper normalization) while allowing a wide range of profiles of varying steepness. (Recall that $\gamma$ and $\beta$ are the negative inner and outer logarithmic slopes of the light profile (respectively), where a Plummer profile corresponds to $\alpha,\>\beta,\>\gamma=2,5,0).$ Lastly, we sample the mass of the stellar component with a flat prior with 25% scatter around $4.3\times 10^{7}\>\text{M}_{\odot},$ as in 2025arXiv250418617T, assuming a constant mass-to-light ratio and that the stellar mass distribution has the same radial density profile as the tracer stars. This is usually a good approximation provided that the tracer profile is complete and that there is no meaningful mass segregation or observational bias in the photometric selection. The latter is typically a good approximation for nearby dwarf galaxies, which have relaxation times longer than a Hubble time due to the presence of dark matter (e.g. 2022MNRAS.510.3531B). If this is not the case, as in star clusters, one needs to include the presence of additional components of varying segregation (e.g. see 2023MNRAS.522.5320D; 2025A&A...693A.104B) in the mass model.
@@
The full implementation of our likelihood modeling with the priors and data we use, and the GravSphere2 code itself, is made publicly available in the Introduction (Sec. 1).
We also note that our priors are generally very broad and have little effect on our results. Some, such as imposing isotropy in the center and barring large negative anisotropies, which we do for the Gaia Challenge sample following 2021MNRAS.505.5686C, are physically motivated constraints. We show explicitly in Appendix B that the space of mass and tracer density models spanned by our priors is much broader than those favored by the data, demonstrating that our priors are sufficiently broad so as to have little effect on our results. For completeness, we have also included marginalized parameter constraints (corner plots) of each of the 19 parameters for all our models in a supplementary repository: https://github.com/dadams42/GS2-results.
@@
All fits were performed using the nested sampling package dynesty (2020MNRAS.493.3132S; 2022zndo...6609296K). This is based on nested sampling techniques (2004AIPC..735..395S; 10.1214/06-BA127) and the ‘dynamic nested sampling technique’ (cite-key) (which we use here). dynesty has been optimized for the inference of posterior distributions (as opposed to evidence estimation), using the bounding method from 2009MNRAS.398.1601F. The sampling method used is based on 10.1214/aos/1056562461; 2015MNRAS.450L..61H; 2015MNRAS.453.4384H. For further information, see: https://dynesty.readthedocs.io/en/latest/index.h
tml
@@
. An example with the specific implementation used for GravSphere2 can be found in the publicly available code release. This sampling method has been shown to outperform Markov-chain Monte Carlo Methods (MCMC), particularly when dealing with complex, high dimensional, and multi-modal posterior spaces (e.g. 2020MNRAS.493.3132S). This is consistent with our own experiments comparing with previous versions of GravSphere that used emcee instead (Foreman_Mackey_2013). Following 2025A&A...693A.104B, we have thus opted to use this improved sampling technique for GravSphere2.